# DESY/PUBDB-2024-01826

### v1

@@ 7 Outlook
There is also significant room for hyperparameter tuning. The used batch size of 4000 is comparatively large, with typical values being less than 1024. The large batch size was chosen for training stability but has been shown to potentially lead to poorer generalisation. The chosen number of training jets of $O(10^{6})$ can be considered a rough lower bound given the number of parameters in the network $\sim 10^{6}$. A natural next step would be to train the network on a much larger number of jets. Further improvements in the network architecture are likely, though this was not explored in the context of these studies.
Subdividing jet flavours into categories with unique signatures, such as $b$-jets into those that decay hadronically and semi-leptonically, or $g\to b\bar{b}$ splittings that do not resemble the typical radiation pattern of a gluon jet, is likely to improve discrimination performance. Additional categories could likewise be included for anti-quarks, which would be helpful in discriminating dijet events where a quark-antiquark pair is expected, such as in $Z\to s\bar{s}$ decays. More generally, much could be gained from event-level tagging, particulary for $s$ quark jets, where discrimination comes primarily from a hard Kaon. Tagging an entire event could require not only a hard Kaon in one jet, but a hard Kaon of the opposite flavour in the other, thus discriminating against Kaons produced during the dressing of a light quark.
The updated design of the IDEA detector concept has the innermost layer of the vertex detector at 1.3 mm instead of 1.7 mm. It will improve the impact parameter resolution and, consequently, the displaced vertex resolutions, thus enhancing the performance of heavy flavour tagging. Further improvement is expected from an ultra-light ALICE ITS3-like vertex detector Freitag:2851362 . An updated version of CLD bacchetta2019cld is being developed with a dedicated RICH PID detector, ARC, which is expected to aid in strange tagging.
@@
The similar performance in Higgsstrahlung events suggests the opportunity to measure the Yukawa coupling of the $s$ quark, and the decent gluon discrimination, especially against heavy quarks, will make gluon final states accessible as well. The much larger $Z$ boson cross-section will also provide opportunities for calibration and performance validation on data before the Higgs boson decay to $s$ quarks is examined, which is likely to reduce experimental uncertainties.
@@ 8 Conclusion
Deep learning techniques have shown great potential in analysing complex jet structures and extracting subtle flavour signatures in jet flavour identification. The transformer-based model presented in this work can be trained considerably more quickly compared to the state-of-the-art graph neural network-based taggers, making it uniquely suited for prospective studies of the developing detector concepts. The discrimination power of this framework called DeepJetTransformer is presented for FCC-ee, allowing the classification of all jet flavours in $e^{+}e^{-}$ collisions at the $Z$ resonance. It should be noted that even though this study focuses on FCC-ee and the IDEA detector, the conclusions are general, and DeepJetTransformer can also be utilised at other collider projects with appropriate adjustments.
A tagging efficiency for $b$-jets of about $99\%$ can be achieved against $s$, $u$, and $d$ jets at a background efficiency of $0.1\%$, pointing to an excellent $b$-jet discrimination, dominantly owing to the secondary vertex reconstruction coming from the expected excellent detector resolution. A $c$-jet tagging efficiency of about $90\%$($70\%$) can be achieved when discriminating from $b$-jets, at a background efficiency of $10\%$($1\%$).
Excellent discrimination can be achieved for $s$-quark tagging against the $b$- and $c$-quark jet background. Against the most challenging background of light jets, a $40\%$ efficiency can be achieved at a background efficiency of $10\%$. Some discrimination can be achieved even between $u$- and $d$-jets. The performance is partially attributed to the inclusion of V0s. Another significant performance enhancement is seen when $K^{\pm}/\pi^{\pm}$ discrimination is included.
It is shown that $Z\to s\bar{s}$ can be efficiently isolated from other hadronic decays of the $Z$ boson. These results show that modern jet flavour tagging techniques can isolate very pure samples of light quark decays originating from vector bosons. We hope that light quark jet tagging will create opportunities for a new category of potential studies at future lepton colliders, including assessment of the feasibility of completely new or more precise measurements and enhancement of the sensitivity to new physics phenomena.
@@ Acknowledgments
We want to thank our CMS colleagues at the IIHE in Brussels, especially A.R. Sahasransu and Lode Vanhecke for their preparatory work, and Emil Bols for their valuable discussions regarding DeepJetTransformer. We would also like to thank Kyle Cormier at the UZH for helpful discussions regarding the feature importance studies.
@@
KG implemented and AI tested the vertex reconstruction algorithm. ADM designed and implemented the DeepJetTransformer architecture. EP adopted the model for the FCC-ee environment and trained and evaluated the performance for several scenarios. KG assessed the classifier performance by isolating $Z\rightarrow s\bar{s}$ events from the exclusive $Z\rightarrow q\bar{q}$ decays. FB and AI supervised and reviewed the work throughout the study. FB, ADM, KG, AI, and EP contributed to the writing and editing of this paper.

### v4

@@ 8 Outlook
There is also significant room for hyperparameter tuning. The used batch size of 4000 is comparatively large, with typical values being less than 1024. The large batch size was chosen for training stability but has been shown to potentially lead to poorer generalisation. The chosen number of training jets of $O(10^{6})$ can be considered a rough lower bound given the number of parameters in the network $\sim 10^{6}$. A natural next step would be to train the network on a much larger number of jets. Further improvements in the network architecture, including adjustments to layer parameters and network structure, are likely possible, though this was not explored in the context of these studies.
Subdividing jet flavours into categories with unique signatures, such as $b$-jets into those that decay hadronically and semi-leptonically, or $g\to b\bar{b}$ splittings that do not resemble the typical radiation pattern of a gluon jet, is likely to improve discrimination performance. Additional categories could likewise be included for anti-quarks, which would be helpful in discriminating dijet events where a quark-antiquark pair is expected, such as in $Z\to s\bar{s}$ decays. More generally, much could be gained from event-level tagging, particularly for $s$ quark jets, where discrimination comes primarily from a high-momentum Kaon. Tagging an entire event could require not only a high-momentum Kaon in one jet but also an oppositely-charged high-momentum Kaon in the other, thus discriminating against Kaons produced during the dressing of a $u$ or $d$ quark, which will not have an oppositely-charged high-momentum Kaon in the other hemisphere.
The updated design of the IDEA detector concept has the innermost layer of the vertex detector at 1.2 mm instead of 1.7 mm. It will improve the impact parameter resolution and, consequently, the displaced vertex resolutions, thus enhancing the performance of heavy flavour tagging. Further improvement is expected from an ultra-light ALICE ITS3-like vertex detector Freitag:2851362 . An updated version of CLD bacchetta2019cld is being developed with a dedicated compact-RICH PID detector, ARC, which is expected to aid in strange tagging.
@@
The similar performance in Higgsstrahlung events suggests the opportunity to measure the Yukawa coupling of the $s$ quark, as attempted in Ref. DuarteCampderros2020 ; Albert:2022mpk , and the decent gluon discrimination, especially against heavy quarks, will make gluon final states accessible as well. The much larger $Z$ boson cross-section will also provide opportunities for calibration and performance validation on data before the Higgs boson decay to $s$ quarks is examined, which is likely to reduce experimental uncertainties.
@@ 9 Conclusion
Deep learning techniques have demonstrated excellent performance in analysing complex jet structures and extracting subtle flavour signatures in jet flavour identification. The short training time of DeepJetTransformer makes it uniquely suited for prospective studies of the developing detector concepts. It should be noted that even though this study focuses on FCC-ee and the IDEA detector, the conclusions are general, and DeepJetTransformer can also be utilised at other collider projects with appropriate adjustments, such as tuning to different detector geometries, jet clustering algorithms, or energy regimes.
These results show that modern jet flavour tagging techniques can isolate very pure samples of $s$ quark decays originating from vector bosons. We hope that strange jet tagging will create opportunities for a new category of potential studies at future lepton colliders, including assessment of the feasibility of completely new or more precise measurements and enhancement of the sensitivity to new physics phenomena.
@@ Acknowledgments
We want to thank our CMS colleagues at the IIHE in Brussels, especially A.R. Sahasransu and Lode Vanhecke for their preparatory work, and Emil Bols for valuable discussions regarding DeepJetTransformer. We would also like to thank Kyle Cormier at the UZH for helpful discussions regarding the feature importance studies. We are grateful to Frank Gaede, Loukas Gouskos, and Michele Selvaggi for their feedback on the manuscript.
@@
KG implemented and AI tested the vertex reconstruction algorithm. ADM designed and implemented the DeepJetTransformer architecture. EP adopted the model for the FCC-ee environment, performed input feature selection, and trained and evaluated the performance for several scenarios. KG assessed the classifier performance by isolating $Z\rightarrow s\bar{s}$ events from the exclusive $Z\rightarrow q\bar{q}$ decays. FB and AI supervised and reviewed the work throughout the study. FB, ADM, KG, AI, and EP contributed to the writing and editing of this paper.